OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - pi/3), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (anonymous):

@shubhamsrg ?

OpenStudy (anonymous):

Help!

OpenStudy (anonymous):

@help123please. @Yating @Tati_Lee @kelliegirl33

OpenStudy (noelgreco):

First, distribute the 2 across the expression enclosed in parentheses.

OpenStudy (anonymous):

Is my explanation correct? "The 5 could be ignored because we are finding the domain. Since each period has an interval of 2pi, the minimum possible domains would be 2pi and 0, or [0, 2pi]. When the first period is 0, 2(x-pi/3)=0, x=pi/3. pi/3+2pi=7pi/3. Therefore, the Xmin and Xmax would be [pi/3,7pi/3]."

OpenStudy (zehanz):

In the formula it says ...cos2(x... It is the 2 that counts here. In general, the period of \(\cos ax\) is \(2\pi/a\). So the period of your function is \(2\pi/2=\pi\). If you want to draw two complete periods, you can set Xmin=0 and Xmax=2pi

OpenStudy (noelgreco):

That's right. The 3 can be ignored as well because it increases the amplitude. Each period, by the way, is pi, not 2pi.

OpenStudy (zehanz):

It is not important that the graph is horizontally and vertically shifted. 2pi is two periods. If you would set Xmin=200 and Xmax=200+2pi, you'd still get two complete periods

OpenStudy (zehanz):

Just think about Ymin and Ymax as well, if you want to be sure the graph is visible :)

OpenStudy (anonymous):

Thank you! Is the final answer of pi/3 and 7pi/3 correct though?

OpenStudy (noelgreco):

yes.

OpenStudy (zehanz):

Think about it: 7pi/3 - pi/3 = 6pi/3=2pi, so, yes it's correct!

OpenStudy (zehanz):

Here is what I described. Look at the values of Xmin and Xmax.